# QVGen: Pushing the Limit of Quantized Video Generative Models

### v1

@@ 5.3 Ablation Studies
Table 3: Ablation results of each component. “Naive” denotes naive QAT in a KD-based manner.
{tabu}
[t!]l|llllll Method Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
$+\Phi$ 63.41 54.75 77.89 15.51 22.98
$+$rank-decay  63.08-0.33 54.67-0.08 77.78-0.11 15.32-0.19 23.01+0.03
To demonstrate the effect of each design, we employ W4A4 Wan 1.3B [64] with 5 dimensions of VBench [25] for ablation studies. More ablations can be found in the appendix.
Effect of different components. We evaluate the contribution of each component in Tab. 5.3. The auxiliary module $\Phi$ (Sec. 4.1) yields substantial performance improvements across all metrics. Furthermore, the rank-decay schedule (Sec. 4.2) effectively eliminates inference overhead, while inducing less than a $0.6\%$ drop in most metrics. It even leads to a slight improvement in Overall Consistency.
Choice of the shrinking ratio $\lambda$. To determine a proper shrinking ratio $\lambda$ in Eq. (14), we conduct experiments in Tab. 5.3. By maintaining the same number of training iterations for each decay phase  [^4] We also employ a fixed total number of training epochs., a small ratio results in an excessively rapid descent of $u$ in Eq. (14) from $1$ to 00, potentially destabilizing the training process. On the other hand, a larger ratio may cause the premature removal of high-contributing components during each phase. An extreme scenario would involve a $100\%$ ratio (i.e., $\lambda=1$), in which the entire ${\mathbf{W}}_{\Phi}$ is removed in a single decay phase, leading to significant performance degradation. Thus, we adopt a moderate ratio of $\lambda=\frac{1}{2}$ in QVGen.
@@
{tabu}
[t!]c|llllll $\quad\quad\lambda\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
$1/2$ 63.08+0.06 54.67+0.05 77.78-0.13 15.32+0.14 23.01+0.12
$3/4$
62.89 54.62 77.91 15.04 22.89
$1$ 61.05 52.48 76.48 13.82 21.81
Table 5: Results of different initial ranks $r$ for Eq. (15). $r=0$ represents “Naive” in Tab. 5.3. We employ $r=32$ in this work.
{tabu}
[t!]c|llllll $\quad\quad r\quad\quad$ Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
@@
Choice of the initial rank $r$. We further present the results for different initial ranks $r$ of ${\mathbf{W}}_{\Phi}$ in Tab. 5.3. As $r$ increases, the performance gains diminish and eventually deteriorate (i.e., at $r=64$). We attribute this trend to the same issue associated with a small shrinking ratio discussed earlier. Specifically, increasing $r$ to $2r$ introduces an additional decay phase, which shortens the training time allocated to each phase and may lead to overly rapid decay (e.g., at $r=64$).
Table 6: Results of different decay strategies. Details of each method can be found in the appendix. “Rank-based” denotes the rank-decay strategy in this work.
{tabu}
[t!]c|lllll Decay Strategy Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Dynamic Degree $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
Sparse-based 61.15 54.06 74.24 13.86 22.52
Res. Q.-based  61.72 54.01 72.41 14.17 22.31
Rank-based 63.08+1.36 54.67+0.61 77.78+3.54 15.32+1.15 23.01+0.49
Analysis of different fine-grained decay strategies. To further demonstrate the effectiveness of our rank-decay strategy, we evaluate alternative decay strategies in Tab. 5.3. Inspired by network pruning [16, 15], we introduce a “Sparse-based” strategy that progressively prunes the largest values in ${\mathbf{W}}_{\Phi}$ during training. Additionally, motivated by residual quantization [38], we design a “Res. Q.-based” strategy, which first quantizes ${\mathbf{W}}_{\Phi}$ into 4×4$\times$4 ×4-bit tensors with the same shape and then progressively removes them one by one. Among all these methods, the “Rank-based” strategy outperforms others across all the metrics by a large margin. This result highlights the superiority of our rank-decay schedule.
Additionally, both “Sparse-based” and “Res. Q.-based” strategies require at least 1.8×1.8$\times$1.8 × training hours at the same setups compared with our “Rank-based” approach (see the appendix).
@@ 5.4 Efficiency Discussion
(a) Model Size (GB)$\downarrow$
(b) Inference Latency (s)$\downarrow$
Figure 6: Evaluation of memory compression and inference acceleration. Blue color and yellow color denote Wan [64] 1.3B and 14B, respectively.
We report the per-step latency of W4A4 DiT components on one $A800$ GPU in Fig. 6 (b), using the same settings as Sec. 5.2. Adapted from the fast CUDA kernel implementation by Ashkboos et al. [1], W4A4 QVGen achieves 1.21×1.21$\times$1.21 × and 1.44×1.$4$ speedups for Wan 1.3B and 14B, respectively. Besides, it exhibits 4$\sim$ 4$\times$ 4 × memory savings compared to the BF16 format, as shown in Fig. 6 (a). Nevertheless, we believe that the acceleration ratio could be further improved with advanced kernel-fusion techniques, which we leave to future work. In addition, it is worth noting that our QVGen adheres to standard uniform quantization, enabling drop-in deployment with existing W4A4 kernels across various devices.

### v3

@@ 4.3 Ablation Studies
To demonstrate the effect of each design, we employ W4A4 Wan 1.31.3B with VBench (Huang et al., 2024b) for ablation studies. More ablations can be found in Sec. J.
Table 3: Ablation results of each component. “Naive” denotes naive QAT in a KD-based manner. “Rank” signifies our rank-decay schedule.
Method
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
$\Phi$ 63.41 54.75 77.89 15.51 22.98
++Rank 63.08 54.67 77.78 15.32 23.01
Effect of different components. We evaluate the contribution of each component in Tab. 3. The auxiliary module $\Phi$ (Sec. 3.1) yields substantial performance improvements across all metrics. Further, the rank-decay schedule (Sec. 3.2) effectively eliminates extra inference overhead, while inducing less than a $0.6\%$ drop in most metrics. It even leads to improvement in Overall Consistency.
Choice of the shrinking ratio $\lambda$. To determine a proper shrinking ratio $\lambda$ in Eq. (13), we conduct experiments in Tab. 5. By maintaining the same training iterations for each decay phase  [^5] We also employ a fixed total number of training epochs., a small ratio results in an excessively rapid descent of $u$ in Eq. (13) from $1$ to 0, potentially destabilizing the training process. On the other hand, a larger ratio may cause the premature removal of high-contributing components during each phase. An extreme scenario would involve a ratio $100\%$ (i.e., $\lambda=1$), in which all ${\mathbf{W}}_{\Phi}$ is removed in a single decay phase, leading to huge performance drops.
@@
$\quad\quad\lambda\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
$1/2$ 63.08 54.67 77.78 15.32 23.01
$3/4$ 62.89 54.62 77.91 15.04 22.89
11 61.05 52.48 76.48 13.82 21.81
Table 5: Results of different initial ranks $r$ for Eq. (14). $r=0$ represents “Naive” in Tab. 3. We employ $r=32$ in this work.
$\quad\quad r\quad\quad$
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
@@
Choice of the initial rank $r$. We further present the results for different initial ranks $r$ of ${\mathbf{W}}_{\Phi}$ in Tab. 5. As $r$ increases, the performance gains diminish and eventually deteriorate (i.e., at $r=64$). We attribute this trend to the same issue associated with a small shrinking ratio discussed earlier. Specifically, increasing $r$ to $2r$ introduces an additional decay phase, which shortens the training time allocated to each phase and may lead to overly rapid decay (e.g., at $r=64$).
Table 6: Results of different decay strategies. Details of each method can be found in Sec. F. “Rank” denotes the rank-decay strategy in this work.
Decay
Strategy
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Dynamic
Degree
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
Sparse 61.15 54.06 74.24 13.86 22.52
Res. Q. 61.72 54.01 72.41 14.17 22.31
Rank 63.08 54.67 77.78 15.32 23.01
Analysis of different fine-grained decay strategies. To further demonstrate the effectiveness of our rank-decay strategy, we evaluate alternative decay strategies in Tab. 6. Inspired by network pruning (Han et al., 2015; 2016), we introduce a “Sparse” strategy that progressively prunes the largest values in ${\mathbf{W}}_{\Phi}$ during training. Additionally, motivated by residual quantization (Li et al., 2017), we design a “Res. Q.” strategy, which first quantizes ${\mathbf{W}}_{\Phi}$ into $\times44$-bit tensors with the same shape and then progressively removes them one by one. Among all these methods, the “Rank” strategy outperforms others across all the metrics by a large margin. Additionally, both “Sparse” and “Res. Q.” strategies require at least $1.8\times$ training hours at the same setups compared with our “Rank” approach (see Sec. F).
@@ 4.4 Efficiency Discussion
[…]
Figure 6: Evaluation of memory compression and inference acceleration. Blue color and yellow color denote Wan 1.31.3B and 1414B, respectively.
We report the per-step latency of W4A4 DiT components on one $A800$ GPU in Fig. 6 (b). Adapted from the CUDA kernel implementation by Ashkboos et al. (2024), W4A4 QVGen achieves $1.21\times$ and $1.44\times$ speedups for Wan 1.31.3B and 1414B, respectively. Besides, it exhibits $\sim 4\times$ memory savings compared to the BF16 format, as shown in Fig. 6 (a). Nevertheless, we believe that the acceleration ratio could be further improved with advanced kernel-fusion techniques and optimization for specific tensor shapes in these models, which we leave to future work. In addition, it is worth noting that our QVGen adheres to standard uniform quantization, enabling drop-in deployment with existing W4A4 kernels across various devices.
Although previous research (Zhang et al., 2025b) mentioned that the current $3D$ full-attention video occupies significant computation in video generation, our QVGen is orthogonal to works that focus on accelerating $3D$ attention and can deliver notable speedups to models that already employ those techniques. For example, when SVG (Xi et al., 2025) is paired with W4A4 QVGen, the model runs $1.70\times$ and $2.63\times$ faster, compared with $1.31\times$ and $1.68\times$ for SVG alone, as shown in Fig. 6 (c).
5 Related Work
Model quantization. Quantization (Jacob et al., 2017) is a predominant technique for minimizing storage and accelerating inference. It can be categorized into post-training quantization (PTQ) (Nagel et al., 2020) and quantization-aware training (QAT) (van Baalen et al., 2020). PTQ compresses models without re-training, making it fast and data-efficient. Nevertheless, it may result in suboptimal performance, especially under ultra-low bit-width (e.g., 33/44-bit). Conversely, QAT applies quantization during training or finetuning and typically achieves higher compression rates with less performance degradation. For DMs, previous quantization research (Li et al., 2023a; He et al., 2023; Huang et al., 2024a; 2025; So et al., 2023; Wu et al., 2024a; Shang et al., 2023; Wang et al., 2024) primarily focuses on image generation. Video DMs, which incorporate complex temporal and spatial modeling, are still challenging for low‑bit quantization. QVD (Tian et al., 2024) and Q-DiT (Chen et al., 2024b) first apply PTQ to convolution-based (Xu et al., 2023b; Guo et al., 2024) and DiT-based (Zheng et al., 2024) video DMs, respectively. Furthermore, ViDiT‑Q (Zhao et al., 2025a) employs mixed‑precision and fine‑grained PTQ to improve performance. However, it experiences video quality loss with 8‑bit activation quantization and has not been extended to more advanced models (Yang et al., 2025; WanTeam et al., 2025). Consequently, QAT for advanced video DMs is urgently needed.
In this work, we identify and address the ineffective low-bit QAT for the video DM. Our proposed method significantly enhances model performance and incurs zero inference overhead in 33/44-bit settings. Moreover, our framework is orthogonal to existing QAT methods, which target gradient estimation (Gong et al., 2019), oscillation reduction (Nagel et al., 2022), etc. We provide more related work about diffusion models in Sec. A.